# Private Heterogeneous Federated Learning Without a Trusted Server Revisited: Error-Optimal and Communication-Efficient Algorithms for Convex Losses

### v1

@@ 3.1 Error-Optimal and Communication-Efficient ISRL-DP FL for Non-Smooth Losses
Our first algorithm is based on the technique of Nesterov smoothing Nesterov (2005): For a non-smooth function $f$, Moreau-Yosida regularization is used to approximate $f$ by the $\beta$-smooth $\beta$-Moreau envelope
@@
We then optimize this smooth function using our Algorithm 1. See Appendix 10 for details.
@@
Theorem 3.1 gives the first optimal rate for non-smooth heterogeneous FL under ISRL-DP. Note that the current best result for heterogeneous FL in Lowy and Razaviyayn (2023a) is suboptimal and holds only for the smooth case. One can combine the same smoothing technique above with the algorithm in Lowy and Razaviyayn (2023a) to obtain a risk bound of $O(1/\sqrt{nN}+(\sqrt{d\ln(1/\delta)}/{\varepsilon n\sqrt{N}})^{2/3})$ for the non-smooth case, which is again suboptimal.
@@
The communication complexity bound in Theorem 3.1 improves over the previous state-of-the-art result for an algorithm achieving optimal excess risk Lowy and Razaviyayn (2023a); recall (3). Moreover, the result of Lowy and Razaviyayn (2023a) assumed i.i.d. silo data, whereas our result holds for the non-i.i.d. case.
@@ 3.2 Error-Optimal and Computationally Efficient ISRL-DP FL for Non-Smooth Losses
We propose another variation of Algorithm 1 that uses subgradients to handle the non-smooth case in a computationally efficient way: see Algorithm 4. Algorithm 4 follows the same structure as Algorithm 1: we iteratively solve a carefully chosen sequence of regularized ERM problems with a ISRL-DP solver and use localization. Compared to Algorithm 1 for the smooth case, Algorithm 4 does not use an accelerated solver (due to non-smoothness). Instead, we use ISRL-DP Minibatch Subgradient method (Algorithm 3) to solve the non-smooth strongly convex ERM problem in each phase of Algorithm 4. There are two key differences between our subroutine Algorithm 3 and the ISRL-DP MB-SGD of Lowy and Razaviyayn (2023a): (i) Instead of the gradient, a subgradient of the non-smooth objective is used in Line 6; (ii) A different and simpler averaging step in Line 11 is used for strongly convex non-smooth losses.
@@
The above subgradient complexity improves over the state-of-the-art gradient complexity Lowy and Razaviyayn (2023a) for an ISRL-DP FL algorithm with optimal excess risk. Lowy and Razaviyayn (2023a) apply Nesterov smoothing to ISRL-DP MB-SGD. As discussed earlier, implementing the smoothing approach is computationally costly. Moreover, the results in Lowy and Razaviyayn (2023a) assume i.i.d. silo data. The subgradient complexity result in Theorem 3.4 also improves over the gradient complexity of Algorithm 1 with smoothing.
An alternative approach is to use the convolutional smoothing technique Kulkarni et al. (2021) to optimize non-smooth functions. For certain regimes, this approach improves the gradient complexity over the subgradient method, with a worse communication complexity compared to the Nesterov smoothing approach. See Appendix 9 for details.
@@ 4 Numerical Experiments
We validate our theoretical findings with numerical experiments on MNIST data. As shown in Figures 3 and 3, our algorithm consistently outperforms Lowy et al. (2023a). We use a similar experimental setup to Lowy et al. (2023a), as outlined below.
@@
We compare our algorithm against the One-pass ISRL-DP MB-SGD of Lowy et al. (2023a). Recall that Lowy et al. (2023a) did not provide theoretical guarantees for their multi-pass ISRL-DP MB-SGD with heterogeneous silos.
@@ 5 Concluding Remarks and Open Questions
We have studied private federated learning in the absence of a trusted server. We characterized the minimax optimal excess risk bounds for heterogeneous ISRL-DP FL, answering an open question posed by Lowy and Razaviyayn (2023a). Further, our algorithms advanced the state-of-the-art in terms of communication and computational efficiency. For smooth losses, the communication complexity of our optimal algorithm matches the non-private lower bound.

### v2

@@ 3.1 Error-Optimal and Communication-Efficient ISRL-DP FL for Non-Smooth Losses
Our first algorithm is based on the technique of Nesterov smoothing (Nesterov, 2005): For a non-smooth function $f$, Moreau-Yosida regularization is used to approximate $f$ by the $\beta$-smooth $\beta$-Moreau envelope
@@
We then optimize this smooth function using our Algorithm 1. See Appendix E for details.
@@
Theorem 3.1 gives the first optimal rate for non-smooth heterogeneous FL under ISRL-DP. Note that the current best result for heterogeneous FL in Lowy and Razaviyayn (2023b) is suboptimal and holds only for the smooth case. One can combine the same smoothing technique above with the algorithm in Lowy and Razaviyayn (2023b) to obtain a risk bound of $O(1/\sqrt{nN}+(\sqrt{d\ln(1/\delta)}/{\varepsilon n\sqrt{N}})^{2/3})$ for the non-smooth case, which is again suboptimal.
@@
The communication complexity bound in Theorem 3.1 improves over the previous state-of-the-art result for an algorithm achieving optimal excess risk (Lowy and Razaviyayn, 2023b); recall (3). Moreover, the result of Lowy and Razaviyayn (2023b) assumed i.i.d. silo data, whereas our result holds for the non-i.i.d. case.
@@ 3.2 Error-Optimal and Computationally Efficient ISRL-DP FL for Non-Smooth Losses
We propose another variation of Algorithm 1 that uses subgradients to handle the non-smooth case in a computationally efficient way: see Algorithm 4. Algorithm 4 follows the same structure as Algorithm 1: we iteratively solve a carefully chosen sequence of regularized ERM problems with a ISRL-DP solver and use localization. Compared to Algorithm 1 for the smooth case, Algorithm 4 does not use an accelerated solver (due to non-smoothness). Instead, we use ISRL-DP Minibatch Subgradient method (Algorithm 3) to solve the non-smooth strongly convex ERM problem in each phase of Algorithm 4. There are two key differences between our subroutine Algorithm 3 and the ISRL-DP MB-SGD of Lowy and Razaviyayn (2023b): (i) Instead of the gradient, a subgradient of the non-smooth objective is used in 6; (ii) A different and simpler averaging step in 11 is used for strongly convex non-smooth losses.
@@
The above subgradient complexity improves over the state-of-the-art gradient complexity (Lowy and Razaviyayn, 2023b) for an ISRL-DP FL algorithm with optimal excess risk. Lowy and Razaviyayn (2023b) apply Nesterov smoothing to ISRL-DP MB-SGD. As discussed earlier, implementing the smoothing approach is computationally costly. Moreover, the results in Lowy and Razaviyayn (2023b) assume i.i.d. silo data. The subgradient complexity result in Theorem 3.4 also improves over the gradient complexity of Algorithm 1 with smoothing.
An alternative approach is to use the convolutional smoothing technique (Kulkarni et al., 2021) to optimize non-smooth functions. For certain regimes, this approach improves the gradient complexity over the subgradient method, with a worse communication complexity compared to the Nesterov smoothing approach. See Appendix D for details.
@@ 4 Numerical Experiments
We validate our theoretical findings with numerical experiments on MNIST data. As shown in Figures 3 and 3, our algorithm consistently outperforms Lowy and Razaviyayn (2023b). We use a similar experimental setup to Lowy and Razaviyayn (2023b), as outlined below.
@@
We compare our algorithm against the One-pass ISRL-DP MB-SGD of Lowy and Razaviyayn (2023b). Recall that Lowy and Razaviyayn (2023b) did not provide theoretical guarantees for their multi-pass ISRL-DP MB-SGD with heterogeneous silos.
@@ 5 Concluding Remarks and Open Questions
We have studied private federated learning in the absence of a trusted server. We characterized the minimax optimal excess risk bounds for heterogeneous ISRL-DP FL, answering an open question posed by Lowy and Razaviyayn (2023b). Further, our algorithms advanced the state-of-the-art in terms of communication and computational efficiency. For smooth losses, the communication complexity of our optimal algorithm matches the non-private lower bound.